OpenStudy (anonymous):

Differentiate the given function: lnx^3 should the next step be 3lnx^2

OpenStudy (apples):

Well, note that here you would use the chain function. The outer function is log(x) and the inner function is x^3. Noting that \[\frac{d}{dx} \left(f(g(x))\right) = f'(g(x)) \cdot g'(x)\] then you'll want to differentiate log(x) and plug x^3 into it, then multiply that by the derivative of x^3.

OpenStudy (anonymous):

good answer apples

OpenStudy (amistre64):

ln(x) dresses down to 1/x: 3ln(x)^2 ------- x

OpenStudy (apples):

That's not correct. You're right that d/dx log(x) = 1/x, however this yields the result \[\frac{1}{x^3} \cdot 3x^2\] which simplifies to \[\frac{3}{x}\]

OpenStudy (anonymous):

if you mean (lnx)^3 then amistre is right if you mean ln(x^3) then apples is right

OpenStudy (amistre64):

...... oh the humanity!! ;)